OpenStudy (anonymous):

write an equation of a line that contains the given point and is perpendicular to the given line. (6,4); y=3x-2

OpenStudy (anonymous):

y=3x+14

OpenStudy (anonymous):

x+3y=18

OpenStudy (anonymous):

perpendicular means opposite reciprocal slope so m=-1/3x y=-1/3x+B (4)=-1/3*(6) +B B=6 y=-1/3x + 6 is the answer